Can you please explain to me how to work this problem?
If 8 workers can paint the walls of a gymnasium in 6 hours, how fast can 12 painters do it?
This is what i tried, and my mom even tried to help me, but we could not figure it out.
8/6= 0.75 so 12(0.75) must give me the correct answer...
Well i got 9 hours which i know isn't correct because if 8 painters can do it in 6 hours, 12 painters should be able to do it faster... right?
Thank you.

Found 2 solutions by richwmiller, checkley77:
Answer by richwmiller(17219)

Before we actually solve it let's look at what you tried.
8/6 does not equal .75 but 4/3 or 1 1/3
6/8=.75
and you are right 12 painters should do it faster than 6 hours unless the new painters are all supervisors.
Now to solve it.
8 painters worked for 6 hours
That's 48 manhours were needed
The number of manhours is the same but we are going to divide 48 by 12 workers. not just 8
48/12=4
It should take the 12 painters 4 hours if they have their own equipment and don't get in each others way.
6*8=x*12
48=x*12
4=x
what is 4? hours worked by 12 workers
what is 6? hours worked by 8 workers




Answer by checkley77(12844)

6*8=48 MANHOURS ARE REQUIRED TO PAINT THE GYM.
48/12=4 HOURS FOR THE 12 WORKERS TO PAINT THE GYM.